{3x+y=5
{2x-y=10
Can you please show all work.

Answer by jb007mi6(53)

it helps to label your equations. lets say (3x+y=5 is "1" and the other is "2"
solve each equation for y
"1" ----- y=-3x+5
"2" ----- y=2x-10
since both equations both equal "y" you can set them equal to eachother
-3x+5=2x-10
solve this equation for "X"
-5x=-15
x=3
put the value of "3" into the first (or second) equation
in "1" 3(3)+y=5
9+y=5
y=-4
check your answers in the other equation as well